Billy is twice as old as his son. In 10 years, the ratio of their age will be 3:5. find their age.
Answer by checkley77(12844)

(x+10)/(2x+10)=3/5
3(2x+10)=5(x+10)
6x+30=5x+50
6x-5x=50-30
x=20 the son'e age now.
2*20=40 dads age now.
Proof:
(20+10)/(40+10)=3/5
30/50=3/5
3/5=3/5